# Modelling the interaction between ethnicity and infectious disease transmission dynamics

### v1

@@ Abstract
During the COVID-19 pandemic, Aotearoa followed an elimination strategy followed by a mitigation strategy, which saw high success and kept health impact low. However, there were inequities in health outcomes, notably that Māori and Pacific Peoples had lower vaccine coverage and experienced higher age-standardised rates of hospitalisation and death. Models provide predictions of disease spread and the burden of disease, which can effectively inform policy, but are often not good at including inequities/heterogeneity. Despite the differences in health outcomes by ethnicity, most models have not explicitly considered ethnic heterogeneities as factors. We developed such a model to investigate the first Omicron wave of the COVID-19 pandemic in Aotearoa, which was the first time there was widespread community transmission of SARS-CoV-2. We analysed three models for contact patterns within and between ethnicities—proportionate mixing, assortative mixing, and empirically derived mixing—and fit them using ethnicity-specific data on reported cases and spatially disaggregated population counts. We found that Māori, Pacific, and Asian contact rates were between 1.17-2.46, 1.60-3.89, and 0.83-0.92 times the European rates, respectively. We then found that from the parameters considered in the model, the disparity in transmission rates explained the majority of the observed disparity in attack rates, while assortativity and vaccine rates explained comparably less.
@@
mathematics; mathematical modelling; epidemiology; health equity; COVID-19; compartmental model;
@@ 1 Introduction
When the COVID-19 pandemic began in 2020, it was predicted that Māori would be disproportionately affected [1], with one model estimating that the COVID-19 infection fatality rates would be 50% higher for Māori compared to non-Māori [2]. A combination of living conditions, lack of access to quality health care, and existing health conditions led to Māori having a higher risk of infection and developing more serious symptoms of COVID-19 [3]. During the first Omicron wave, Māori and Pacific Peoples had the highest per capita rates of cases and hospitalisations, with odds ratios of hospitalisation of 2.03 and 1.75 for Māori and Pacific Peoples, respectively [4]. This was consistent with historical inequities in infectious disease burden experienced by Māori and Pacific Peoples [5, 6, 7, 8, 9, 10].
@@
During the COVID-19 pandemic, Aotearoa changed its policy approach to COVID-19 at multiple stages. Aotearoa initially responded to the spread of COVID-19 with an elimination strategy. The elimination strategy aimed to completely eliminate COVID-19 from the country. This approach worked and resulted in about 30 cases and 0.4 deaths per hundred thousand population by August 2020 [11]. When vaccines became widely available, a mitigation strategy was used instead [12]. Starting in February, 2021, COVID-19 vaccines began to be administered to border and managed isolation and quarantine workers [13] and then the public thereafter. From March to May 2021, there was a large gap in the proportion of vaccinated Māori and Pacific Peoples compared to other ethnicities. At this time, the government implemented the mitigation strategy via the "traffic light" system formally known as the COVID-19 Protection Framework. This framework implemented social distancing requirements, mask mandates, and COVID-19 vaccination certificates and their requirements [14]. From May 2020 until the Omicron waves started in January 2022, there was very little community transmission. Starting in 2022, when there was high vaccination coverage [15], there was some transmission of Omicron cases from isolation facilities at the border where international arrivals were being isolated. This resulted in a major wave that peaked on February 25th, 2022, with 24,226 daily confirmed cases. After this, the borders were reopened, and international travel was allowed in March 2022. Starting in July, an influx of immune-evasive variants was observed through the border. Most of these cases did not result in detected community transmission; a small amount did, resulting in large outbreaks [16]. Although the first and second doses of the vaccine had reduced efficacy against the Omicron variant compared to previous variants, it was shown that the booster doses were effective at stopping severe disease [17].
@@
Models add a lot of value to predictions of the spread and burden of disease and can effectively inform policy for prevention measures but are often not very good at including inequities/heterogeneity. Without these considerations, the models cannot address or predict these issues as effectively. Previous models have often struggled to account for ethnic heterogeneities despite their effect on disease spread [18], which has limited their ability to assess health equity. There are many models that use socioeconomic status [19] or age [20, 21, 22] as factors, but models that consider ethnicity as a factor [23, 24] are rarer. The mathematical models that were used to inform the Aotearoa policy approach to COVID-19 lacked the capability to adequately answer questions about health inequities and policy effectiveness. This limitation was identified as a major gap in the literature in these policy informing works [25, 26]. As such, there is a need to develop models that can answer these questions and consider a greater number of heterogeneities.
@@
A method developed by Ma et al. fit considered ethnicity explicitly using data from Long Island and New York City during late April, 2020 [23]. In this paper, they used a compartmental model and split the population into five ethnicity subgroups: non-Hispanic white, Hispanic or Latino, non-Hispanic Black/African American, non-Hispanic Asian, and Multiracial/Other. Contact patterns were estimated using an ethnicity-specific contact rate, which was the average number of people interacted with each day, and an assortativity constant, $\epsilon$ , which measured the extent to which people are more likely to contact other people of the same ethnicity. Their model was first fit to a cross-sectional serosurvey from New York in late April 2020, using a proportionate mixing assumption ($\epsilon=0$), where people met according to population sizes and ethnicity-specific contact rates with no preference for interacting with any ethnicity. From this fit it was found that minorities, such as Hispanics or Latinos and non-Hispanic Black/African Americans, had higher contact rates compared to non-Hispanic white people, with contact rates in New York City being 2.25, 1.62, 0.86, and 1.28 times the non-Hispanic white rate for Hispanic or Latino, non-Hispanic Black/African American, non-Hispanic Asians, and Multiracial/Other, respectively.
@@
Ma et al. then estimated a value for the assortativity constant based on the extent to which people of the same ethnicity tend to be clustered together in the same census blocks, with an assumption of density-based proportionate mixing within each block. This process gave more weight to the ethnic distribution of higher population blocks. The assortativity constant was chosen to minimise the difference between a census-informed matrix and a matrix formed using the total populations and the assortativity constant. The result of this fit was that 46% and 39% of an ethnicity’s contacts were purely intra-ethnic contact in New York and Long Island, respectively; the disparity in contact rates was lowered to 1.62, 1.35, 0.90, and 1.17 times the non-Hispanic rate in New York City for Hispanic or Latino, non-Hispanic Black/African American, non-Hispanic Asians, and Multiracial/Other, respectively, but the disparity in rates was still present.
@@
Here, we applied a similar method to Ma et al. and aimed to estimate ethnicity-specific contact patterns by fitting a model to data on confirmed cases of COVID-19 during Aotearoa’s first Omicron wave in 2022. We also aimed to try and understand the relative contributions of different drivers (vaccination rates, contact rates, and assortative contact patterns) to observed trends. We made two main adjustments to the Ma et al. method: we considered a vaccinated population in addition to an unvaccinated population and we changed the way contacts were estimated within census blocks, which allowed us to work directly with the census inform matrix. We then applied our model to the first Omicron wave in Aotearoa to analyse differences in transmission rate by ethnicity. These transmission rates were then used as parameters to investigate various scenarios to analyse the spread of COVID-19 in the model.
@@ 2 Methods
$\frac{d\text{S}}{dt}$ $=-\beta\text{SI}$ (1)
$\frac{d\text{E}}{dt}$ $=\beta\text{SI}-\sigma\text{E}$ (2)
$\frac{d\text{I}}{dt}$ $=\sigma\text{E}-\gamma\text{I}$ (3)
$\frac{d\text{R}}{dt}$ $=\gamma\text{I}$ (4)
where S, E, I, and R refer to the susceptible, exposed, infectious, and recovered populations, respectively. We set the mean latent period, $1/\sigma$, to 3 days and the mean infectious period, $1/\gamma$, to 4 days to match the Ma et al. study. There is some evidence that the Omicron variant had a shorter latent and infectious period than the wild-type variant [27]. However, as we are mainly interested in cumulative attack rates here, our results are not sensitive to choices for these parameters. The per capita transmission rate was defined in terms of the basic reproduction number, $R_{0}$, and the total population, $N$, $\beta=\gamma R_{0}/N$.
We then extended this model in the same way as Ma et al. to incorporate multiple ethnic groups by considering a per capita transmission matrix, $\boldsymbol{\beta}$, which governs the rate at which transmission happens within and between the considered ethnic groups. The per capita transmission matrix elements, $\beta_{ij}$, were the average proportion of susceptible people in ethnic group $i$ an infectious individual from ethnic group $j$ infected each day. In addition to this (beyond the Ma et al. extension), we also considered a susceptible vaccinated group, Sv, and redefined S to be the unvaccinated susceptible group.
$\frac{d\textbf{S}}{dt}$ $=-(\boldsymbol{\beta}\textbf{I})\circ\textbf{S}$ (5)
$\frac{d\textbf{S}_{v}}{dt}$ $=-(1-v_{e})(\boldsymbol{\beta}\textbf{I})\circ\textbf{S}_{v}$ (6)
$\frac{d\textbf{E}}{dt}$ $=(\boldsymbol{\beta}\textbf{I})\circ\textbf{S}+(1-v_{e})(\boldsymbol{\beta}\textbf{I})\circ\textbf{S}_{v}-\sigma\textbf{E}$ (7)
@@
$\frac{d\textbf{R}}{dt}$ $=\gamma\textbf{I}$ (9)
In these equations, $v_{e}$ refers to the vaccine effectiveness, which we set to be $0.3$, as it has been shown that the vaccines had reduced effectiveness against infection via Omicron [28]; $\circ$ denotes element-wise multiplication; and S, $\textbf{S}_{v}$, E, I, and R denote column vectors containing the compartmental populations for each group. The ethnic groups considered were Māori, Pacific Peoples, Asian people, and European/Others, as they are the major ethnic groups within Aotearoa and are consistently identified in health and population data and were identified in our data [15, 29, 30, 31].
@@ 2.1 Transmission matrix
We defined the per capita contact matrix using a similar method to Ma et al. [23]. We considered that on average, each ethnic group interacted with a different number of people per day; this was modelled with an ethnicity-specific contact rate, $\hat{a}_{k}$, where $k$ denoted the ethnicity. We assumed that the proportion $q$ of contacts resulting in infection was the same for all ethnicities. We define the transmission vector as $\textbf{a}=q\hat{\textbf{a}}$. The transmission vector was the average number of people infected per day by an infectious person in a fully susceptible population. The differences in transmission rates between ethnicities are interpreted as being caused by differences in contact rates, i.e. social factors, not biological factors. Using this definition, $q$ and $\hat{a}_{k}$ were not separately identifiable, and so we expressed the model in terms of $\boldsymbol{a}$ and aimed to estimate $\boldsymbol{a}$ by fitting the model to data as described below.

### v2

@@ Abstract
During the COVID-19 pandemic, Aotearoa followed an elimination strategy followed by a mitigation strategy, which saw high success and kept health impact low. However, there were inequities in health outcomes, notably that Māori and Pacific Peoples had lower vaccine coverage and experienced higher age-standardised rates of hospitalisation and death. Models provide predictions of disease spread and burden, which can effectively inform policy, but are often less good at including inequities/heterogeneity. Despite the differences in health outcomes, most models have not explicitly considered ethnic heterogeneities as factors. We developed such a model to investigate the first Omicron wave of the COVID-19 pandemic in Aotearoa, which was the first widespread community transmission of SARS-CoV-2. We analysed three models for contact patterns within and between ethnicities: proportionate, assortative, and unconstrained mixing, which were fit using ethnicity-specific data on reported cases and spatially disaggregated population counts. We found that Māori, Pacific, and Asian transmission rates were between 1.08-2.46, 1.50-3.89, and 0.80-0.92 times the European rates, respectively. We then found that from the parameters considered in the model, the disparity in ethnic transmission rates explained the majority of the observed ethnic disparity in attack rates, while assortativity and vaccination rates explained comparatively less.
@@
mathematical modelling; epidemiology; health equity; COVID-19; compartmental model;
1 Plain language summary
Mathematical models of how an epidemic spreads through a population are important to inform policy decisions about public health interventions. Despite the relevance to disease spread, many models do not include important variables such as socioeconomic status and ethnicity. Here we investigate the first Omicron wave of COVID-19 in Aotearoa New Zealand with a model that splits the population into 10 groups relating to an individual’s disease and vaccination status: susceptible, susceptible, exposed, infectious, and recovered and unvaccinated, vaccinated, or vaccinated and booster. We further split the population into 4 ethnicity groups, corresponding to the main ethnicity classifications in Aotearoa: Māori, Pacific, Asian, and European/Other. We used this model to estimate how much contact rates would need to differ between ethnicities in order to explain observed differences in infection rates. We did this with three different assumptions about how many contacts occur with and between groups: (1) people mix with all ethnicities equally; (2) people are more likely to mix with others in the same ethnicity group; and (3) people are more likely to mix with others living in the same geographical area. From this, we estimated that Māori, Pacific, and Asian contact rates were 1.08-2.46, 1.50-3.89, and 0.80-0.92 times the European rates, respectively. We then found that the disparity in ethnic contact rates was sufficient to explain the majority of the observed disparity in infection rates by ethnicity, while preference for interacting with your own ethnic group and differences in vaccination rates between ethnicities explained comparatively less of the observed disparity.
@@ 2 Introduction
When the COVID-19 pandemic began in 2020, it was predicted that Māori would be disproportionately affected [jones_why_2020], with one model estimating that the COVID-19 infection fatality rates would be 50% higher for Māori compared to non-Māori [steyn_estimated_2020]. A combination of living conditions, lack of access to quality health care, and existing health conditions led to Māori having a higher risk of infection and developing more serious symptoms of COVID-19 [king_covid-19_2020]. During the first Omicron wave, Māori and Pacific Peoples had the highest per capita rates of cases and hospitalisations, with odds ratios of hospitalisation of 2.03 and 1.75 for Māori and Pacific Peoples, respectively [whitehead_inequities_2023]. This was consistent with historical inequities in infectious disease burden experienced by Māori and Pacific Peoples [rice_remembering_2019, wilson_differential_2012, esr_measles_2019, chapple_death_2018, bennett_rising_2021, cheung_severe_2024].
@@
During the COVID-19 pandemic, Aotearoa changed its policy approach to COVID-19 at multiple stages. Aotearoa initially responded to the spread of COVID-19 with an elimination strategy. The elimination strategy aimed to completely eliminate COVID-19 from the country. This approach worked and resulted in about 30 cases and 0.4 deaths per hundred thousand population by August 2020 [baker_successful_2020]. When vaccines became widely available, a mitigation strategy was used instead [ministry_of_health_nz_managing_2024]. Starting in February, 2021, COVID-19 vaccines began to be administered to border and managed isolation and quarantine workers [ardern_first_2021] and then the public thereafter. From March to May 2021, there was a large gap in the proportion of vaccinated Māori and Pacific Peoples compared to other ethnicities. At this time, the government implemented the mitigation strategy via the “traffic light” system formally known as the COVID-19 Protection Framework. This framework implemented social distancing requirements, mask mandates, and COVID-19 vaccination certificates and their requirements [parliamentary_counsel_office_covid-19_2021]. From May 2020 until the Omicron waves started in January 2022, there was very little community transmission. Starting in 2022, when there was high vaccination coverage [ministry_of_health_new_2023], there was some transmission of Omicron cases from isolation facilities at the border where international arrivals were being isolated. This resulted in a major wave that peaked on February 25th, 2022, with 24,226 daily confirmed cases. After this, the borders were reopened, and international travel was allowed in March 2022. Starting in July, an influx of immune-evasive variants was observed through the border. Most of these cases did not result in detected community transmission; a small amount did, resulting in large outbreaks [douglas_tracing_2022]. Although the first and second doses of the vaccine had reduced efficacy against the Omicron variant compared to previous variants, it was shown that the booster doses were effective at stopping severe disease [bowen_omicron_2022].
@@
Models add a lot of value to predictions of the spread and burden of disease and can effectively inform policy for prevention measures but are often not very good at including inequities/heterogeneity. Without these considerations, the models cannot address or predict these issues as effectively. Previous models have often struggled to account for ethnic heterogeneities despite their effect on disease spread [williamson_opensafely_2020], which has limited their ability to assess health equity. There are many models that use socioeconomic status [goodfellow_covid-19_2024] or age [lyra_covid-19_2020, miller_disease_2020, davies_age-dependent_2020] as factors, but models that consider ethnicity as a factor [ma_modeling_2021, lee_ethnic_2023] are rarer. The mathematical models that were used to inform the Aotearoa policy approach to COVID-19 lacked the capability to adequately answer questions about health inequities and policy effectiveness. This limitation was identified as a major gap in the literature in these policy informing works [datta_modelling_2023, rosenstrom_covsim_2024]. As such, there is a need to develop models that can answer these questions and consider a greater number of heterogeneities.
@@
A method developed by Ma et al. considered ethnicity explicitly using data from Long Island and New York City during late April, 2020 [ma_modeling_2021]. In this paper, they used a compartmental model and split the population into five ethnicity subgroups: non-Hispanic white, Hispanic or Latino, non-Hispanic Black, non-Hispanic Asian, and Multiracial/Other. Contact patterns were estimated using an ethnicity-specific contact rate, which was the average number of people interacted with each day, and an assortativity constant, $\epsilon$ , which measured the extent to which people are more likely to contact other people of the same ethnicity. Their model was first fit to data from a cross-sectional seroprevalence study from New York in late April 2020, using a proportionate mixing assumption ($\epsilon=0$), where people met according to population sizes and ethnicity-specific contact rates with no preference for interacting with any ethnicity. From this fit it was found that minorities had higher contact rates compared to non-Hispanic White people, with contact rates in New York City being 2.25, 1.62, 0.86, and 1.28 times the non-Hispanic White rate for the Hispanic or Latino, non-Hispanic Black American, non-Hispanic Asian, and Multiracial/Other groups, respectively.
@@
Ma et al. then estimated a value for the assortativity constant based on the extent to which people of the same ethnicity tend to be clustered together in the same census blocks, with an assumption of density-based proportionate mixing within each block. This process gave more weight to the ethnic distribution of higher population blocks. The assortativity constant was chosen to minimise the difference between a census-informed matrix and a matrix formed using the total populations and the assortativity constant. The result of this fit was that 46% and 39% of an ethnicity’s contacts were purely intra-ethnic contact in New York and Long Island, respectively; the disparity in ethnic contact rates was lowered to 1.62, 1.35, 0.90, and 1.17 times the non-Hispanic rate in New York City for the Hispanic or Latino, non-Hispanic Black, non-Hispanic Asian, and Multiracial/Other groups, respectively, but the disparity in rates was still present.
@@
Here, we applied a similar method to Ma et al. and aimed to estimate ethnicity-specific contact patterns by fitting a model to data on confirmed cases of COVID-19 during Aotearoa’s first Omicron wave in 2022. We also aimed to try and understand the relative contributions of different drivers (vaccination rates, contact rates, and assortative contact patterns) to observed trends. We made two main adjustments to the Ma et al. method: we considered a vaccinated population in addition to an unvaccinated population and we changed the way contacts were estimated within census blocks, which allowed us to work directly with the census-informed matrix. We then applied our model to the first Omicron wave in Aotearoa to analyse differences in transmission rate by ethnicity. These transmission rates were then used as parameters to investigate various scenarios to analyse the spread of COVID-19 in the model.
@@ 3 Methods
$\frac{dS}{dt}$ $=-\beta SI$ (1)
$\frac{dE}{dt}$ $=\beta SI-\sigma E$ (2)
$\frac{dI}{dt}$ $=\sigma E-\gamma I$ (3)
$\frac{dR}{dt}$ $=\gamma I$ (4)
where S, E, I, and R refer to the susceptible, exposed, infectious, and recovered populations, respectively. We set the mean latent period, $1/\sigma$, to 3 days and the mean infectious period, $1/\gamma$, to 4 days to match the Ma et al. study (A table of parameter values can be seen in Table 3.2.3). There is some evidence that the Omicron variant had a shorter latent and infectious period than the wild-type variant [park_inferring_2023]. However, as we are mainly interested in cumulative attack rates here, our results are not sensitive to choices for these parameters. The per capita transmission rate was defined in terms of the basic reproduction number, $R_{0}$, and the total population, $N$, $\beta=\gamma R_{0}/N$.
We then extended this model in the same way as Ma et al. to incorporate multiple ethnic groups by considering a per capita transmission matrix, $\boldsymbol{\beta}$, which governs the rate at which transmission happens within and between the considered ethnic groups. The per capita transmission matrix elements, $\beta_{ij}$, were the average proportion of susceptible people in ethnic group $i$ an infectious individual from ethnic group $j$ infected each day. In addition to this (beyond the Ma et al. extension), we also stratified the susceptible population according to whether they were: (i) unvaccinated ($S$); (ii) had received the primary vaccination course, normally two doses ($S_{V}$); (iii) had received a third vaccine dose, i.e. a booster ($S_{vb}$). There were relatively few people who received one dose and not a second, as such we grouped them into the primary vaccination We created separate compartments to track when these populations become exposed and infectious (E, $\textbf{E}_{v}$, $\textbf{E}_{vb}$, I, $\textbf{I}_{v}$, and $\textbf{I}_{vb}$, respectively). We assumed that the effect of vaccinations was both a reduction in susceptibility and a reduction in transmission when infectious, and that the effect of vaccinations/boosters were constant over our study period. This leads to the following model:
$\frac{d\textbf{S}}{dt}$ $=-\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}$ (5)
$\frac{d\textbf{S}_{v}}{dt}$ $=-v_{a}\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}_{v}$ (6)
$\frac{d\textbf{S}_{vb}}{dt}$ $=-v_{ab}\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}_{vb}$ (7)
$\frac{d\textbf{E}}{dt}$ $=\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}-\sigma\textbf{E}$ (8)
$\frac{d\textbf{E}_{v}}{dt}$ $=v_{a}\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}_{v}-\sigma\textbf{E}_{v}$ (9)
$\frac{d\textbf{E}_{vb}}{dt}$ $=v_{ab}\boldsymbol{\beta}(\textbf{I}+v_{t}\textbf{I}_{v}+v_{tb}\textbf{I}_{vb})\circ\textbf{S}_{vb}-\sigma\textbf{E}_{vb}$ (10)
@@
$\frac{d\textbf{I}_{v}}{dt}$ $=\sigma\textbf{E}_{v}-\gamma\textbf{I}_{v}$ (12)
$\frac{d\textbf{I}_{vb}}{dt}$ $=\sigma\textbf{E}_{vb}-\gamma\textbf{I}_{vb}$ (13)
$\frac{d\textbf{R}}{dt}$ $=\gamma(\textbf{I}+\textbf{I}_{v}+\textbf{I}_{vb})$ (14)
In these equations, $1-v_{a}$ and $1-v_{ab}$ represent vaccine effectiveness against infection for those in the two-dose and three-dose compartments, respectively, while $1-v_{t}$ and $1-v_{tb}$ are the same for vaccine effectiveness against transmission in infected individuals; $\circ$ denotes element-wise multiplication; and S, $\textbf{S}_{v}$, $\textbf{S}_{vb}$, E, $\textbf{E}_{v}$, $\textbf{E}_{vb}$, I, $\textbf{I}_{v}$, $\textbf{I}_{vb}$, and R denote column vectors containing the compartmental populations for each group. The ethnic groups we considered were Māori, Pacific, Asian, and European/Other, as they are the major ethnic groups within Aotearoa and are consistently identified in health and population data and were identified in our data [ministry_of_health_new_2023, statsnz_statistical_2020, te_whatu_ora_covid-19_2025, te_whatu_ora_populations_2023].
Vaccines had reduced effectiveness against infection via Omicron and as such we needed to estimate their effectiveness [cromer_predicting_2023]. Hao et al. [hao_predicting_2025] estimated vaccine effectiveness as function of time since the second vaccine dose or the third vaccine dose. We did not model time-dependent immunity explicitly, but instead chose representative values corresponding to the timing of the vaccine rollout relative to the modelled wave. Since the primary vaccine rollout in Aotearoa occurred mainly in the last few months of 2021 [ministry_of_health_new_2023], we considered that those who had not received a booster doses would have about 120 days of immunity decay. The bulk of the booster rollout occurred in late 2021 to early 2022, just before the modelled wave, so we considered that those in the boosted compartment would have about 50 days of immunity decay. Values for vaccine effectiveness can be seen in Table 3.2.3.
@@ 3.1 Transmission matrix
We defined the per capita contact matrix using a similar method to Ma et al. [ma_modeling_2021]. We considered that on average, each ethnic group interacted with a different number of people per day; this was modelled with an ethnicity-specific contact rate, $\hat{a}_{k}$, where $k$ denoted the ethnicity. We assumed that the proportion $q$ of contacts resulting in infection was the same for all ethnicities. We define the transmission vector as $\textbf{a}=q\hat{\textbf{a}}$. Each component, $a_{k}$, of the transmission vector was the average number of people infected by an infectious individual of ethnicity $k$ in a fully susceptibility population. The differences in transmission rates between ethnicities are interpreted as being caused by differences in contact rates, i.e. social factors, not biological factors. Using this definition, $q$ and $\hat{a}_{k}$ were not separately identifiable, and so we expressed the model in terms of $\boldsymbol{a}$ and aimed to estimate $\boldsymbol{a}$ by fitting the model to data as described below.